OpenStudy (anonymous):

So I've tried this multiple times and can't get any of the given answers... help? ? Washington High School has a total of 90 boys and girls who play sports. If the number of boys is 10 more than three times the number of girls, how many girls play sports at this high school? 20 30 70 80

OpenStudy (anonymous):

B =3G +10 B+G=90 B=90-G G=20

OpenStudy (anonymous):

x = girls 3x + 10 = boys 4x + 10 = 90 4x = 80 x = 20 20 Girls' 70 Boys

OpenStudy (anonymous):

first option 20